OpenStudy (anonymous):

Given a geometric sequence whose sum of the first 10 terms is 4 and whose sum from the 11th to the 30th term is 48, find the sum from the 31st to the 60th term.

OpenStudy (cwrw238):

Sum of 10 = a * (r^4 - 1) -------- = 4 r- 1

OpenStudy (anonymous):

@cwrw238 why is it r^4? I thought it's r^10.

OpenStudy (anonymous):

n is 10 not 4

OpenStudy (amistre64):

hmm, given is: \[S_n=\frac{1-r^n}{1-r}\] \[S_{10}=4=\frac{1-r^{10}}{1-r}\] \[S_{30-10}=48=\frac{1-r^{20}}{1-r}\]

OpenStudy (anonymous):

@hitten101 yes yes :)

OpenStudy (shubhamsrg):

you have been given the sum upto first 10 terms =4 you have also been given the sum upto first 30 terms = 4 + 48 =52 and you have 2 eqns with 2 variables ->solve for a and r now calculate sum for first 60 terms from that subtract sum of first 30 terms.. this should help..

OpenStudy (anonymous):